A glass wedge of angle 0.01 radius is illuminated by a monochromatic light of wavelength 6330 falling normally it. At what distance from the of the the 20th fringe is observed
1
Expert's answer
2020-02-21T10:01:09-0500

Let's clarify the condition. So,


A glass wedge ("\mu=1.5" ) of angle 0.01 radian is illuminated by a monochromatic light of wave length 6330 A falling normally on it. At what distance from the edge of the wedge, will the 20 th fringe be observed by reflected light.


For bright bands


"2\mu d=(2n+1)\frac{\lambda}{2}"


"d=x_n\cdot \tan\alpha"


"2\mu \cdot x_n\cdot \tan \alpha=(2n+1)\frac{\lambda}{2}"


"x_n=\frac{(2n+1)\lambda}{4\mu \alpha}=\frac{(2\cdot 20+1)\cdot 633\cdot10^{-9}}{4\cdot 1.5\cdot 0.01}\approx0.00043 m=0.43 mm"